### v1

# Weyl’s Relations, Integrable Matrix Models and Quantum Computation

Abstract
Starting from a generalization of Weyl’s relations in finite dimension $N$, we show that the Heisenberg commutation relations can be satisfied in a specific $N$-1 dimensional subspace, and display a linear map for projecting operators to this subspace. This setup is used to construct a hierarchy of parameter-dependent commuting matrices in $N$ dimensions. This family of commuting matrices is then related to Type-1 matrices representing quantum integrable models. The commuting matrices find an interesting application in quantum computation, specifically in Grover’s database search problem. Each member of the hierarchy serves as a candidate Hamiltonian for quantum adiabatic evolution and, in some cases, achieves higher fidelity than standard choices — thus offering improved performance.
June 20, 2025
I Introduction
In this work, we present a link between Weyl’s relations, commuting matrices and quantum computation, within the context of finite-dimensional matrices representing quantum systems in finite-dimensional Hilbert spaces. To this end, we construct a class of mutually commuting real symmetric $N\times N$ matrices that depend linearly on a parameter $x$, representing quantum integrable systems in finite dimensions. These are obtained from an algebra involving three matrices (${\cal E},{\cal S},{\bf\Gamma}$), in a structure that is closely related to generalizations of matrices appearing in Weyl’s relations Weyl ; Schwinger ; Schwinger2 ; Tekumalla . Weyl’s concern was with the emergence of the Heisenberg commutation relations in the limit of large dimensions. We display variants of the Weyl matrices defining Weyls’s relations in finite dimensions, which yield canonical commutators in well defined subspaces defined by the orthogonality with respect to a single specific state. Using these, we construct a hierarchy of commuting matrices and find that they coincide with certain linear combinations of Type-1 commuting matrices previously obtained by us in shastry ; shastry1 ; haile ; haile1 ; scaramazza ; scaramazza1 ; yuzbashyan , using entirely different arguments.
We show that the commuting matrices found here generalize a model Hamiltonian studied in the context of quantum computation. It is used to display the quantum advantage over classical methods in the problem of a search in a random database — i.e., Grover’s algorithm as implemented through adiabatic evolution Grover1 ; Grover2 ; Grover-Analysis ; Farhi-Guttman ; QAA_Theory ; QAA ; LAA ; Berkeley ; Childs ; Young .
A detailed review of Type-1 matrices is available in yuzbashyan . These constitute a class of finite-dimensional matrices that aim to capture the essence of quantum integrability at the level of a matrix representation of models, rather than at the more conventional Hamiltonian level involving physical degrees of freedom. Here we derive the Type-1 commuting matrices from the commutation relation (39).
We first discuss the Weyl relations in the large $N$ limit in Section II, where the Weyl matrices $A$ and $B$ [see Eq. (5) and Eq. (3)] are introduced and a review of how the Heisenberg algebra emerges from them is provided taking a careful continuum limit Schwinger ; Tekumalla ; comment-Santhanam . In Section III, we summarize earlier generalizations of the Weyl’s operator basis of Eq. (24), which are used for creating a basis of operators representing maximally entangled states. We then introduce a matrix $C$ Eq. (25), a new generalization of the Weyl matrices, which together with $B$ defines an algebra in Eq. (27) that plays an important role in the subsequent analysis. This is a simplified version of the algebra found in Section IV in Eq. (39), which is finally employed to generate the commuting matrices in Section V. In Section VI, we discuss the use of the Type-1 matrices in implementing the Grover algorithm through adiabatic computation. By studying the fidelity, we show that the higher order matrices lead to an increase in the efficiency of computation. In Section VII, we make some final remarks.
II Weyl matrices and the limit $N\to\infty$
Weyl proposed an important pathway for understanding the Heisenberg algebra
$[Q,P]=i\hbar\mathds{1},$ (1)
emerging from a finite-dimensional setting, in the limit of infinite dimensions. Taking the limit of infinite dimensions is necessary to avoid a familiar contradiction, arrived at by taking the trace of the Heisenberg algebra in finite dimensions. Weyl’s work was followed by others detailing and extending his arguments Weyl ; Schwinger ; Schwinger2 ; Tekumalla . Since our work relies on yet another extension of Weyl’s work, before describing it we first provide a quick summary of the known results.
Weyl begins with a relation for a pair of $N\times N$ unitary matrices $A$ and $B$,
$AB=BAe^{i\omega_{0}},$ (2)
where $\omega_{0}=\frac{2\pi}{N}$ and $e^{i\omega_{0}}$ is the $N^{th}$ root of unity with integer $N$ — taken to infinity at the end. It follows from iteration that $A^{N}=\mathds{1}=B^{N}$. The form of the matrices is determined as follows: first we set $B$ to be a diagonal with diagonal entries $\{b_{0},b_{1},\ldots,b_{N-1}\}$, and $b_{r}=e^{i\omega_{0}r}$ (thus satisfying $B^{N}=\mathds{1}$), i.e.,
$B=\sum_{n=0}^{N-1}b_{n}|n\rangle\langle n|.$ (3)
On taking the matrix elements of Eq. (2), we get (for $10\leq\{i,j\}\leq N-$ 1) the conditions
$A_{ij}b_{j}=e^{i\omega_{0}}\;b_{i}A_{ij}.$ (4)
Together with the choice $b_{n}=e^{i\omega_{0}n}$, this implies that the only non-zero elements of $A$ are of the form $A_{j,j+1}$, and we may represent $A$ as
$A=\sum_{n=0}^{N-1}|n-1\rangle\langle n|,$ (5)
with periodic boundary conditions $n\equiv n+N$. Thus $A$ is the left-shift operator on a ring of sites of length $N$. $A$ is diagonalized in the plane-wave basis
$|k_{r}\rangle=\frac{1}{\sqrt{N}}\sum_{n=0}^{N-1}e^{ik_{r}n}|n\rangle,$ (6)
where $k_{r}=r\,\omega_{0}$ and $10\leq r\leq N-$ 1, so that $A|k_{r}\rangle=e^{ik_{r}}|k_{r}\rangle$, and
$A=\sum_{r}e^{ik_{r}}|k_{r}\rangle\langle k_{r}|.$ (7)
Consider first the limit of $N\to\infty$. Weyl shows that the Heisenberg algebra emerges from the Weyl relation (2) by expanding both sides for large $N$. This process requires a definition of the position operator $\hat{Q}$ and the momentum operator $\hat{P}$, which are defined by
$A=e^{i\xi\hat{P}/\hbar},\quad B=e^{i\eta\hat{Q}},$ (8)
together with the requirement that $\xi,\eta$ are each of ${\cal O}(\frac{1}{\sqrt{N}})$. With this, Eq. (2) vanishes to ${\cal O}(\frac{1}{N^{2}})$, provided
$\xi\eta\,[\hat{Q},\hat{P}]=i\hbar\omega_{0}.$ (9)
This leads to the Heisenberg algebra Eq. (1), provided
$\xi\eta=\omega_{0}.$ (10)
Following Schwinger Schwinger2 , we make a symmetric choice,
$\xi=\eta=\sqrt{\omega_{0}}.$ (11)
We must also deal with the situation that the eigenvalues of $\log B$, namely $i\omega_{0}n$, are discrete with a small spacing, while the eigenvalues of $\hat{Q}\to x$ are quasi-continuous — indeed, we would like to take the continuum limit as $N\to\infty$. This is achieved by the Weyl scaling, where
$x\leftrightarrow\sqrt{\omega_{0}}\;n,\quad\textrm{and}\quad|x\rangle\leftrightarrow\frac{1}{\sqrt[4]{\omega_{0}}}|n\rangle,$ (12)
so that lattice constant $\sqrt{\omega_{0}}$ shrinks as $N$ increases. With
$x_{1}=\sqrt{\omega_{0}}\,n,\quad x_{2}=\sqrt{\omega_{0}}\,m,$ (13)
the Dirac delta overlap of the continuum states is related to the Kronecker delta overlap in the discrete case through the relation
$\langle x_{1}|x_{2}\rangle=\delta(x_{1}-x_{2})\leftrightarrow\frac{1}{\sqrt{\omega_{0}}}\delta_{n,m}.$ (14)
A completely parallel situation exists for the momentum operator $\hat{P}$, which also has a continuous set of eigenvalues in the large $N$ limit case. We skip its discussion since we only need Eq. (11) and Eq. (14) below.
On the other hand, we now regard $N$ as finite and compute the logarithms of the matrices and evaluate their commutators. We begin with
$\log B=i\omega_{0}\sum_{n=0}^{N-1}n\;|n\rangle\langle n|$ (15)
and
$\log A=i\omega_{0}\sum_{r=0}^{N-1}r|k_{r}\rangle\langle k_{r}|$ (16)
Using Eq. (6), we obtain
$\log A$ $=$ $i\pi\left(1-\frac{1}{N}\right)\mathds{1}+i\frac{\pi}{N}\sum_{n\neq m}\frac{|n\rangle\langle m|b_{m}-|m\rangle\langle n|b_{n}}{b_{n}-b_{m}},$ (17)
$=$ $i\pi\left(1-\frac{1}{N}\right)\mathds{1}+i\frac{\pi}{N}\sum_{n\neq m}\left[\cot\frac{\pi}{N}(n-m)-i\right]|n\rangle\langle m|,$ (18)
the second form (18) illustrating the useful result that convolution with the cotangent in real space is equivalent to the first derivative for a class of functions on the lattice Shastry-inverse .
Going further, we compute the matrix element of the commutator comment-Santhanam . Denoting $\Delta=\omega_{0}(n-m)$, and treating $\Delta$ formally as a continuous variable, we obtain the matrix element
$\langle n|\,[\log B,\log A]\,|m\rangle$ $=$ $-\frac{\omega_{0}}{N}{\Delta}\sum_{r=0}^{N-1}re^{i\Delta r}$ (19)
$=$ $i\frac{\omega_{0}}{N}{\Delta}\frac{d}{d\Delta}\left(e^{i\frac{1}{2}\Delta(N-1)}\frac{\sin\frac{1}{2}\Delta N}{\sin\frac{1}{2}\Delta}\right).$
Further, Eq. (13) implies $\Delta=\sqrt{\omega_{0}}(x_{1}-x_{2})$ and using Eq. (11) we rewrite Eq. (19) as
$\sqrt{\omega_{0}}\langle x_{1}|[\log B,\log A]|x_{2}\rangle=i\frac{\omega_{0}}{N}{(x_{1}-x_{2})}\frac{d}{d(x_{1}-x_{2})}\left(e^{ic_{0}(x_{1}-x_{2})}\frac{\sin\sqrt{\frac{1}{2}\pi N}(x_{1}-x_{2})}{\sin\sqrt{\frac{\pi}{2N}}(x_{1}-x_{2})}\right)\!\!,$ (20)
where $c_{0}=\frac{1}{2}\sqrt{\omega_{0}}(N-1)$. By using Dirichlet-Bonnet’s result Dirichlet-Bonnet , we can set
$\lim_{N\to\infty}\frac{\sin\sqrt{\frac{1}{2}\pi N}(x_{1}-x_{2})}{\sin\sqrt{\frac{\pi}{2N}}(x_{1}-x_{2})}\to\frac{2\pi}{\sqrt{\omega_{0}}}\delta(x_{1}-x_{2}).$ (21)
Upon using $(x_{1}-x_{2})\delta^{\prime}(x_{1}-x_{2})=-\delta(x_{1}-x_{2})$, Eq. (20) reduces to
$\sqrt{\omega_{0}}\langle x_{1}|[\log B,\log A]|x_{2}\rangle=-i(\omega_{0})^{\frac{3}{2}}\delta(x_{1}-x_{2}).$ (22)
Using the replacements $\log B\to i\sqrt{\omega_{0}}\hat{Q}$ and $\log A\to i\sqrt{\omega_{0}}\hat{P}/\hbar$ and canceling the common $\omega_{0}^{\frac{3}{2}}$ term, we recover the matrix elements of the Heisenberg algebra (1),
$\langle x_{1}|[\hat{Q},\hat{P}]|x_{2}\rangle=i\hbar\delta(x_{1}-x_{2}).$ (23)
III Weyl matrices with N finite
Generalizations of Weyl’s matrices $A,B$ for finite $N$ have recently received much interest in quantum computation. These are useful as they form a basis of $N^{2}$ operators representing maximally entangled states genweyl1 ; genweyl2 ; genweyl3 , and are referred to as the Weyl operator basis. Each element of the $N^{2}$-dimensional operator basis is explicitly given by
$U_{n,m}=\sum_{r=0}^{N-1}e^{ink_{r}}|k_{r}\rangle\langle k_{r+m}|,$ (24)
where $n$ and $m$ range over $10,N-$ 1. These matrices are extensions of the Weyl matrix $A$ given in the form (7).
We turn to a generalization of Weyl’s operators in a different direction guided by the question: Is it possible to preserve Eq. (1) as an operator identity while keeping the dimension $N$ finite, when acting on a well-defined subspace? One is faced with a similar question when quantizing constrained field theories, where the constraint is satisfied only as an operator acting on the physical states. This leads us to define another class of matrices, where the analog of Eq. (1) is satisfied on a $(N-1)\times(N-1)$ subspace of the original $N\times N$ space of states.
Towards this, we now define a matrix operator $C$ through the formula
$C=\frac{1}{2}\sum_{n\neq m}\frac{|n\rangle\langle m|-|m\rangle\langle n|}{b_{n}-b_{m}}.$ (25)
This matrix is closely related to $\log A$ of Eq. (17), with the omission of the $b_{m}$ and $b_{n}$ in the numerator of the second term. Notice that $C$ in Eq. (25) and $B$ in Eq. (3) are defined in terms of $b_{n}$ and the following property is independent of the choice of $b_{n}$. The commutator of $C$ with $B$ in Eq. (3) is computed as
$[C,B]$ $=$ $\sum_{n\neq m}(|n\rangle\langle m|)$ (26)
$=$ $\mathds{1}-N|\psi\rangle\langle\psi|,$ (27)
where the projection operator is constructed from the normalized “flat state” $|\psi\rangle$, which is defined as
$|\psi\rangle$ $=$ $\frac{1}{\sqrt{N}}\sum_{n}|n\rangle=|k_{r=0}\rangle,$ (28)
and is also the zero wave-vector ($k_{r}=0$) state from Eq. (6). From Eq. (27), we see that
$[C,B]\,|\Psi\rangle=|\Psi\rangle,$ (29)
for any state $|\Psi\rangle=\sum_{r\neq 0}\alpha_{r}|k_{r}\rangle$ for arbitrary $\alpha_{r}$. This implies that Eq. (29) remains valid as long as $|\Psi\rangle$ is contained in the $(N-1)$-dimensional subspace spanned by the basis $|k_{r}\rangle$ with $r\neq 0$ — i.e., the subspace orthogonal to the state $|k_{r=0}\rangle$. The trace of $[C,B]$ vanishes due to the contribution from the last term. In this sense, the conjugate $N\times N$ matrices $C$ and $B$ come close to the canonical infinite-dimensional Heisenberg algebra (1). A particular advantage of the structure of $C$ is that the commutator (27) is valid for an arbitrary choice of $b_{n}$ including real values. Below, we show the close connection of $B$ and $C$ with matrices that arise in the context of the Type-1 integrable family, both using real values for the $b_{n}$.
We note that the matrix $C$ can be written in the $k$ representation by using Eq. (6) as
$C=\frac{N}{2\pi}\sum_{r=0}^{N-1}k_{r}|k_{r}\rangle\langle k_{r+1}|,$ (30)
which is a minimally off-diagonal version of the log-Weyl matrix $\log A$ from Eq. (18). The algebra satisfied by $C$ seems to be quite rich. We quote one further result found by commuting the expressions in Eq. (16) with the one in Eq. (30):
$[\log A,C]=-iC.$ (31)
This result is reminiscent of the Fock commutator of $\left[p,\frac{1}{2}(px+xp)\right]=-i\hbar$, which is used to obtain the virial theorem in quantum theory. The commutator $[\log B,C]$ is not related simply to $C$, so the analogy is only partial. We also note that the matrix $C$ in Eq. (30) is similar to matrix $U_{11}$ of Eq. (24), and is obtained, up to an overall prefactor, when the $e^{ik_{r}}$ is replaced by $ik_{r}$.
In summary, the three matrices: $C$ in Eq. (25) and $A$ and $B$ in Eq. (2), and their logarithms introduced in Eq. (17) and Eq. (15), satisfy the basic commutation results (27) and (31) for the special choice of $b_{n}=e^{i\omega_{0}n}$. These relations, with the exception of Eq. (31), generalize to arbitrary choices of $b_{n}$. We may qualitatively view $B$ as the position variable itself, and then $A$ can be viewed as the discrete translation operator. The role of $C$ is somewhat context driven. In Eq. (31) $C$ is analogous to the Fock operator $xp+px$, but in Eq. (29), C is analogous to the momentum operator. In Eq. (29) we saw that $C$ and $B$ are canonically conjugate in a subspace which excludes the state $|\psi\rangle$ of Eq. (28). Thus, transforming arbitrary operators so that they annihilate the flat state is expected to be very useful — and we follow up on this idea below. We will refer to $C$ and $B$ as generalized Weyl matrices when the entries $b_{n}$ are chosen arbitrarily.
We also mention the work on a truncated harmonic oscillator restricted to operate in an $N$-dimensional Hilbert space Buchdahl ; Buchdahl-further1 ; Buchdahl-further2 , which leads to a similar algebra as Eq. (27), but with fixed expressions for the matrices.
IV The algebra of Type-1 matrices and connection with generalized Weyl matrices
A brief summary of the Type-1 matrix models is provided next, interested readers can find more details in the review yuzbashyan . In shastry ; shastry1 ; haile ; haile1 ; yuzbashyan ; scaramazza , we developed a theory of sets of real symmetric $N\times N$ matrices depending linearly on a real parameter, say $x$, as finite dimensional prototypes of quantum integrable systems. These emerge from a study of the conservation laws of quantum integrable systems, such as the 1-d Heisenberg and 1-d Hubbard models, when restricted to finite sizes. The matrix formulation has the advantage of being “blind” to the originating quantum model, while capturing elements related to their quantum integrability. Members of the commuting set are in the form $\alpha(x)=a+xA$, where $a$ and $A$ are $N\times N$ real symmetric matrices obeying certain constraint relations. A convenient formulation to define this set of matrices is to introduce $N$ objects $Z_{j}$ that play the role of a basis:
$Z_{j}=|j\rangle\langle j|+x\sum_{k\neq j}\frac{1}{\epsilon_{j}-\epsilon_{k}}\left\{\;\gamma_{j}\gamma_{k}(|j\rangle\langle k|+|k\rangle\langle j|)-\gamma_{j}^{2}|k\rangle\langle k|-\gamma_{k}^{2}|j\rangle\langle j|\;\right\},$ (32)
where the commutation $[Z_{j},Z_{l}]=0$ for all $j,l$ is established, with an unconstrained set of $12N+$ 1 parameters $\{\gamma_{j},\epsilon_{j}\}$ and $x$. We may then write $\alpha(x)=\sum_{j}c_{j}Z_{j}$, and varying the coefficients $c_{j}$ generates the family of mutually commuting Type-1 matrices.
In this program of constructing finite-dimensional matrices representing quantum integrable systems, we unexpectedly encountered an algebraic structure scaramazza that bears considerable resemblance to the above problem (27). We show here that our integrable matrix model, and its conservation laws can be obtained purely from this algebra. This calculation thereby offers a new perspective on our model and provides further insights.
Our main cast consists of a set of matrix operators ${\cal S}$, ${\cal E}$, ${\bf\Gamma}$, and ${\bf D}$ with real entries acting on the $N$-dimensional real vector space ${\cal R}_{N}$. The antisymmetric matrix ${\cal S}$ is defined through its matrix elements
${\cal S}_{ij}=x\;(1-\delta_{ij})\frac{\gamma_{i}\gamma_{j}}{\epsilon_{i}-\epsilon_{j}},\;\;\mbox{ as }{\cal S}=\sum_{i,j}|i\rangle\,{\cal S}_{ij}\,\langle j|\;,$ (33)
where $x$ is a real parameter, $\{\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{N}\}$ are $N$ real variables, and $\{\gamma_{1},\gamma_{2},\ldots,\gamma_{N}\}$ are $N$ real numbers normalized by the condition
$\sum_{i=1}^{N}\gamma_{i}^{2}=1.$ (34)
We also need three diagonal matrices:
${\cal E}=\sum_{j=1}^{N}\epsilon_{j}|j\rangle\langle j|,$ (35)
${{\bf D}}=\sum_{j=1}^{N}\gamma_{j}^{2}|j\rangle\langle j|.$ (36)
and a projection operator
${\bf\Gamma}$ $=$ $|\gamma\rangle\langle\gamma|,\;\;\mbox{where}$ (37)
$|\gamma\rangle$ $=$ $\sum_{j=1}^{N}\gamma_{j}|j\rangle.$ (38)
It follows from the normalization of $\{\gamma_{i}\}$ that $\mbox{Tr}\,{{\bf\Gamma}}=1$ and $\mbox{Tr}\,{\bf D}=1$.
It is seen by comparing Eq. (33) with Eq. (25) that ${\cal S}$ is the same as $C$ introduced earlier, when the constants $\gamma_{i}\to 1$, and $\epsilon_{i}\to b_{i}$. With these changes, we see that ${\cal E}$ becomes $B$ of Eq. (3) and $\bf D$ reduces to $\mathds{1}$. Finally, we see that ${\bf\Gamma}$ is proportional to $|\psi\rangle\langle\psi|$ where $|\psi\rangle$ is given in Eq. (28), and the flat states $|\gamma\rangle$ map onto $|\psi\rangle$. With these changes, the algebra discussed below can be viewed as a generalization of Section III.
The commutator of ${\cal E}$ and ${\cal S}$ is easily seen to be
$[{\cal E},{\cal S}]=\,x\,\left({\bf\Gamma}-{{\bf D}}\right),$ (39)
with a vanishing trace of both sides, as expected. In the special case of $\gamma_{i}=\frac{1}{\sqrt{N}}$ for all $i$, we see that Eq. (39) reduces to Eq. (27), and $x$ is analogous to the Planck’s constant.
Below we uncover $N$ mutually commuting operators $I_{n}$ [see Eq. (56)] using the three building blocks ${\cal E},{\bf\Gamma},{\cal S}$. Among these, there is another natural candidate, Eq. (53), that can be viewed as the Hamiltonian. Indeed we will see later that Eq. (53) is a natural Hamiltonian for a Grover type quantum search algorithm, as discussed in Farhi-Guttman ; QAA ; QAA_Theory ; LAA .
IV.a A linear map on operators
Guided by our earlier discussion [see paragraphs below Eq. (31)], we now find a systematic method for mapping any operator to a related one which annihilates the projection operator ${\bf\Gamma}$. These are found by first analyzing the action of arbitrary operators ${\cal Q}$ on the state $|\gamma\rangle$. Guided by the form of Eq. (38), we now introduce a useful decomposition for any symmetric operator ${\cal Q}$
${\cal Q}={\cal Q}_{||}+{\cal Q}_{\perp},$ (40)
where ${\cal Q}=\sum_{ij}|i\rangle Q_{ij}\langle j|$ and $Q_{ij}=Q_{ji}$. The two pieces of ${\cal Q}$ have the property that
${\cal Q}|\gamma\rangle={\cal Q}_{||}|\gamma\rangle\;\mbox{ and }{\cal Q}_{\perp}|\gamma\rangle=0.$ (41)
More explicitly,
${\cal Q}_{||}=\sum_{ij}|i\rangle\langle i|\;Q_{ij}\left(\frac{\gamma_{j}}{\gamma_{i}}\right),$ (42)
and
${\cal Q}_{\perp}=-\frac{1}{2}\sum_{ij}\frac{1}{\gamma_{i}\gamma_{j}}\Bigl{\{}|i\rangle\gamma_{j}-|j\rangle\gamma_{i}\Bigr{\}}\;Q_{ij}\;\Bigl{\{}\gamma_{j}\langle i|-\gamma_{i}\langle j|\Bigr{\}}.$ (43)
It is clear that ${\cal Q}_{\perp}$ annihilates the vector $|\gamma\rangle$, because the terms $\Bigl{\{}\gamma_{j}\langle i|-\gamma_{i}\langle j|\Bigr{\}}$ annihilate the state $|\gamma\rangle$ for every pair $i,j$. Using the symmetry of ${\cal Q}$, Eq. (43) also implies the orthogonality condition
${\cal Q}_{\perp}.{\bf\Gamma}=0={\bf\Gamma}.{\cal Q}_{\perp}.$ (44)
This construction can also be viewed as a linear map ${\cal Q}\to{\cal Q}_{||}$. Towards this end, we define the equal amplitude (flat sum) state $|\Phi\rangle$:
$|\Phi\rangle=\sum_{j}|j\rangle.$ (45)
It is useful to define an invertible diagonal operator
${\tilde{\bm{\gamma}}}$ $=$ $\sum_{j}\gamma_{j}|j\rangle\langle j|,\;\;\;\;\mbox{ so that }$
$|\gamma\rangle$ $=$ ${\tilde{\bm{\gamma}}}|\Phi\rangle.$ (46)
In terms of ${\tilde{\bm{\gamma}}}$ it is easy to see that
${\cal Q}_{||}\equiv\sum_{j}\;\;\langle j|{\tilde{\bm{\gamma}}}^{-1}{\cal Q}{\tilde{\bm{\gamma}}}|\Phi\rangle\;|j\rangle\langle j|.$ (47)
In summary, the linear operator map (47) helps us to decompose an arbitrary real symmetric operator ${\cal Q}$ into two components, ${\cal Q}={\cal Q}_{||}+{\cal Q}_{\perp}$, where the latter annihilates the state $|\gamma\rangle$.
IV.b Analogy with stochastic equations
A simple analogy between the construction in Eqs. (42), (43), and (47), and stochastic time evolution equations in nonequilibrium physics may be helpful here. In the study of master equations, the equilibrium state can be used to perform a similarity transformation that enables the interpretation of the transition operator as a Hamiltonian, and the master equation itself as the Euclidean version of Schrödinger’s equation MasterEquation1 ; MasterEquation2 ; MasterEquation3 . If we consider a state vector $|P\rangle$, with components representing the probabilities of the various basis states, and a transition operator ${\cal T}$ satisfying a master equation
$\partial_{t}|P\rangle=-{\cal T}|P\rangle,\;\;\;\;\langle\Phi|{\cal T}=0,$ (48)
where $|\Phi\rangle$ defined in Eq. (45) gives equal weight to every configuration $|i\rangle$, then the total probability is conserved. We can write the equilibrium state as
$|{P_{eq}}\rangle={{\tilde{\bm{\gamma}}}}^{2}|\Phi\rangle=\sum_{j}\gamma_{j}^{2}|j\rangle,$ (49)
with positive weights for each configuration, so that the system flows to equilibrium provided ${\cal T}|P_{eq}\rangle=0.$
In order to transform this to the Hamiltonian formulation one performs a similarity transformation
$|P\rangle={\tilde{\bm{\gamma}}}|\Psi\rangle,\;\;H={\tilde{\bm{\gamma}}}^{-1}{\cal T}{\tilde{\bm{\gamma}}},\;\;$ (50)
giving rise to the imaginary time Schröedinger equation
$\partial_{t}|\Psi\rangle=-H|\Psi\rangle.\;\;$ (51)
The equilibrium distribution under this similarity transformation gives rise to the ground state wave function $|\Psi_{0}\rangle\equiv{\tilde{\bm{\gamma}}}^{-1}|P_{eq}\rangle={\tilde{\bm{\gamma}}}|\Phi\rangle$, which satisfies the conditions
$H|\Psi_{0}\rangle=0=\langle\Psi_{0}|H.$ (52)
For a class of transition matrices this prescription gives a Hermitian Hamiltonian. The state $|\gamma\rangle$ in Eq. (46) now plays the role of the ground state $|\Psi_{0}\rangle$. This is the route taken in connecting Dyson’s Brownian motion of matrices with the Calogero-Sutherland model of interacting particles in one dimension.
V The parameter-dependent matrix Hamiltonian and its conservation laws
Using the above algebra, we construct a Hamiltonian $H$ as
$H={\cal E}+x({\bf\Gamma}-\mathds{1})\;\equiv I_{1}$ (53)
We also refer to $H$ as $I_{1}$, since, as we show below, it belongs to a family of matrices,
$I_{m}={\cal E}^{m}+[{\cal E}^{m},{\cal S}]_{\perp},$ (54)
that commute among each other.
From Eq. (39) we note that the commutator $[{\cal E}^{m},{\cal S}]$ is linear in $x$. We thus define
$[{\cal E}^{m},{\cal S}]_{\bot}\equiv x\,K_{m},$ (55)
so that the the conservation laws (54) now read
$I_{m}={\cal E}^{m}+x\ K_{m},$ (56)
with $m=1,2,\ldots N$. Here the effective Planck’s constant $x$ is arbitrary, but common to all the members of this commuting family. Similar constants arise in most quantum integrable systems, e.g., the $U$ parameter in the 1-d Hubbard model. These satisfy for all $n,m\leq N$
$[H,I_{m}]=0,$ (57)
$~{}[I_{n},I_{m}]=0.$ (58)
In fact, we will begin by showing that the first member of the sequence in Eq. (56) is the Hamiltonian (53). For this purpose we use Eq. (39) to write
$[{\cal E},{\cal S}]_{\perp}=x({\bf\Gamma}-{{\bf D}})_{\perp}=x({\bf\Gamma}-\mathds{1}),$ (59)
where we used ${{\bf D}}_{\perp}=0$, a result valid for any diagonal operator, thereby proving that $I_{1}=H$. It may be useful to record the explicit form of the next conservation law
$I_{2}={\cal E}^{2}+x\left({\cal E}{\bf\Gamma}+{\bf\Gamma}{\cal E}-{\cal E}-\mathds{1}\;\langle\gamma|{\cal E}|\gamma\rangle\right).$ (60)
Let us now prove Eq. (57) for general $m$. The commutator is quadratic in $x$, hence we need the vanishing of the three terms $x^{\alpha}$, with $\alpha=0,1,2$ separately. The constant term is trivial since $[{\cal E},{\cal E}^{m}]=0$. The $O(x)$ term requires
$[{\cal E}^{m},[{\cal E},{\cal S}]_{\bot}]=[{\cal E},[{\cal E}^{m},{\cal S}]_{\bot}].$ (61)
We note that the perpendicularity condition can be dropped on both sides, since the parallel part commutes with ${\cal E}$ and ${\cal E}^{m}$. We invoke the Jacobi identity
$[{\cal E}^{m},[{\cal E},{\cal S}]]+[{\cal S},[{\cal E}^{m},{\cal E}]]+[{\cal E},[{\cal S},{\cal E}^{m}]]=0,$ (62)
and since the middle term is identically zero, this implies Eq. (61). The $O(x^{2})$ term requires
${\bf\Gamma}.[{\cal E}^{m},{\cal S}]_{\perp}=[{\cal E}^{m},{\cal S}]_{\perp}.{\bf\Gamma}.$ (63)
Now both sides vanish by the definition (44) of the orthogonality, and thereby we have proved Eq. (57) for any integer $m$.
V.a Mutual Commutation of $I_{n}$
If we assume that the spectrum of $H$ is nondegenerate, the commutation of the operators (58) is implied by a standard argument from linear algebra. However it is worthwhile proving Eq. (58) directly without taking recourse to making assumptions about the spectrum of $H$. Note that we have proved Eq. (57) purely algebraically, without enquiring into the form of ${\cal S}$ and ${{\bf D}}$. It might be possible to do so for Eq. (58) as well, but we now present a proof that uses the explicit solution for ${\cal S}$ and ${{\bf D}}$.
Returning to Eq. (54) and using the explicit expressions (33), (35), and (36), we first work out
$\frac{1}{x}\;[{\cal E}^{m},S]_{ij}=\gamma_{i}\gamma_{j}\frac{\epsilon_{i}^{m}-\epsilon_{j}^{m}}{\epsilon_{i}-\epsilon_{j}}.$ (64)
It is useful to introduce the vectors with circular brackets
$|p)=\sum_{j}\gamma_{j}\epsilon_{j}^{p}|j\rangle,$ (65)
so that $|0)=|\gamma\rangle$. We further define the overlap and projector ${\cal J}$
$(p|q)$ $=$ $\sigma_{p+q},\;\;{\cal J}(p,q)=|p)(q|,$
$\mbox{where}\;\sigma_{p}$ $=$ $\sum\gamma_{i}^{2}\epsilon_{i}^{p}.$ (66)
Thanks to the normalization, we have $\sigma_{0}=1$. In terms of these we find
$\frac{1}{x}\;[{\cal E}^{m},S]=\sum_{r=0}^{m-1}{\cal J}(m-1-r,r)-m\sum_{j}\gamma_{j}^{2}\epsilon_{j}^{m-1}|j\rangle\langle j|.$ (67)
We can construct the perpendicular parts of these operators immediately from the definitions: In terms of a set of diagonal operators
${\cal E}^{p}=\sum_{j}\epsilon_{j}^{p}|j\rangle\langle j|$ (68)
we obtain
$\frac{1}{x}\;[{\cal E}^{m},S]_{\bot}=K_{m}=\sum_{r=0}^{m-1}\left({\cal J}(m-1-r,r)-{\cal E}^{m-1-r}\;\sigma_{r}\right),$ (69)
so that
$I_{m}={\cal E}^{m}+x\ \sum_{r=0}^{m-1}\left({\cal J}(m-1-r,r)-{\cal E}^{m-1-r}\;\sigma_{r}\right).$ (70)
With this explicit form for $I_{m}$, we compute the commutator
$[I_{m},I_{n}]=0,$ (71)
the $O(x)$ term vanishes since we can ignore the $\bot$ constraint under the commutator with diagonal operators:
$[{\cal E}^{m},[{\cal E}^{n},S]_{\bot}]-[{\cal E}^{n},[{\cal E}^{m},S]_{\bot}]=[{\cal E}^{m},[{\cal E}^{n},S]]-[{\cal E}^{n},[{\cal E}^{m},S]].$ (72)
The right-hand side vanishes upon applying the Jacobi identity. Thus the only non trivial condition to check is the $O(x^{2})$ term $[K_{m},K_{n}]=0$, so we write
$[K_{m},K_{n}]$ $=$ $\sum_{r=0}^{m-1}\sum_{s=0}^{n-1}[\left({\cal J}(m-1-r,r)-{\cal E}^{m-1-r}\;\sigma_{r}\right),\left({\cal J}(n-1-s,s)-{\cal E}^{n-1-s}\;\sigma_{s}\right)],$ (73)
$=$ $\sum_{r=0}^{m-1}\sum_{s=0}^{n-1}\left(\sigma_{\mu}\ \bar{{\cal J}}(r,s)+\sigma_{r}\ \bar{{\cal J}}(s,\mu)+\sigma_{s}\ \bar{{\cal J}}(\mu,r)\right),$
where we abbreviated $\bar{{\cal J}}(a,b)={\cal J}(a,b)-{\cal J}(b,a)$, $\mu=m+n-2-r-s$, and used ${\cal E}^{p}|q)=|p+q)$. The term vanishes upon performing the sum, as we next show.
V.b Proof of cancellation of Eq. (73):
Let us write $m\to m+1$ and $n\to n+1$ for convenience and write the (1×)(-1$\times$)( - 1 × ) right-hand side of Eq. (73) in the form
${\cal Y}=\sum_{r=0}^{m}\sum_{s=0}^{n}\{s,r|m+n-r-s\},$ (74)
[…]
$\{a,b|c\}$ $=$ $[a,b;c]+[c,a;b]+[b,c;a],$
$~{}[a,b;c]$ $=$ $\bar{{\cal J}}(a,b)\sigma_{c}.$ (75)
In fact, the detailed form of $[a,b;c]$ is not important. The only property needed is the skew symmetry
$~{}[a,b;c]=-[b,a;c].$ (76)
From these definitions it is easily seen that
$\{a,b|c\}$ $=$ $\{c,a|b\}=\{b,c|a\},\;\;\;\;\mathrm{(i)}$
$\{a,b|c\}$ $=$ $-\{b,a|c\},\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\mathrm{(ii)}$
$\{a,a|c\}$ $=$ $0.\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\mathrm{(iii)}$ (77)
We now separate ${\cal Y}$ into two parts assuming $m>n$,
${\cal Y}$ $=$ ${\cal Y}_{1}+{\cal Y}_{2},$
${\cal Y}_{1}$ $=$ $\sum_{r=0}^{n}\sum_{s=0}^{n}\{s,r|m+n-r-s\},$
${\cal Y}_{2}$ $=$ $\sum_{r=n+1}^{m}\sum_{s=0}^{n}\{s,r|m+n-r-s\}.$ (78)
We first note that ${\cal Y}_{1}$ vanishes since the ranges of $s$ and $r$ are identical, so we can switch them, $r\leftrightarrow s$, and use property (ii) in Eq. (77), so that ${\cal Y}_{1}=-{\cal Y}_{1}=0$. To analyze ${\cal Y}_{2}$, we fix $r$ in its new range, and observe that the resulting range of the integer $(m+n-r-s)$ coincides with that of $s$. For a given $r$ we can now see that as $s$ varies, there are two cases: (a) the integer $(m+n-r-s)$ equals $s$, which vanishes by using property (iii) in Eq. (77), or case (b) integer $(m+n-r-s)$ is distinct from $s$, in which case we can exchange these, $(m+n-r-s)\leftrightarrow s$, and then use property (ii) in Eq. (77) to show that the sum of these terms vanishes.
We have thus shown that the algebra (39) directly leads to the conservation laws $I_{n}$. We can take the linear sums $\sum_{j}c_{j}I_{j}$ and these are still constants of motion. The earlier construction of the Type-1 family can be related to the $I_{n}$ straightforwardly. Indeed, we can express the operators $I_{m}$ of Eq. (56) in terms of $Z_{j}$ given by Eq. (32) as
$I_{n}=\sum_{j}\epsilon_{j}^{n}Z_{j}.$ (79)
Note that this relationship implies that only $N-1$ of the infinite sequence of $I_{n}$ are linearly independent.
V.c Spectra, recurrence relation, general formula, and $I_{n}$ as a polynomial in $H$:
The spectrum of a general type-1 matrix was determined in haile . The eigenvalues $\xi_{\alpha}$ and unnormalized eigenvectors $|\varphi_{\alpha}\rangle$ (shared by all matrices in the commuting family) are
$\xi_{\alpha}=x\sum_{i=1}^{N}\frac{d_{i}\gamma_{i}^{2}}{\lambda_{\alpha}-\epsilon_{i}},\quad\varphi^{(i)}_{\alpha}=\frac{\gamma_{i}}{\lambda_{\alpha}-\epsilon_{i}},\quad\alpha=1,\dots,N.$ (80)
where $d_{i}$ are the eigenvalues of the $x=0$ part of the matrix and $\lambda_{\alpha}$ are the $N$ roots of the equation
$\sum_{i=1}^{N}\frac{\gamma_{i}^{2}}{\lambda_{\alpha}-\epsilon_{i}}=\frac{1}{x}.$ (81)
For $I_{n}$ we have $d_{i}=\epsilon_{i}^{n}$, i.e., the eigenvalues of $I_{n}$ read
$\eta^{(n)}_{\alpha}=x\sum_{i=1}^{N}\frac{\epsilon_{i}^{n}\gamma_{i}^{2}}{\lambda_{\alpha}-\epsilon_{i}}.$ (82)
Note that when $x\to 0$, $\lambda_{\alpha}\to\epsilon_{i}$ and $\eta^{(n)}_{\alpha}\to\epsilon_{i}^{n}$. When $x\to\infty$, one $\lambda_{\alpha}$ (say $\lambda_{N}$) diverges, while others remain finite. Eq. (81) implies
$\lambda_{N}\to x\sum_{i}\gamma_{i}^{2}=x.$ (83)
The corresponding eigenvector (upon normalization) $|\varphi_{N}\rangle\to|\gamma\rangle$ and the eigenvalue $\eta^{(n)}_{N}\to 0$. Indeed, Eqs. (41) and (55) imply that $|\gamma\rangle$ is an eigenvector of $I_{n}$ in $x\to\infty$ limit with eigenvalue 0.
Eigenvalues of $I_{n}$ are polynomials in $\lambda_{\alpha}$ of order $n$. Indeed, substituting the identity
$\epsilon_{i}^{n}=(\epsilon_{i}-\lambda_{\alpha})(\epsilon_{i}^{n-1}+\epsilon_{i}^{n-2}\lambda_{\alpha}+\dots+\epsilon_{i}\lambda^{n-1}_{\alpha}+\lambda^{n-1}_{\alpha})+\lambda^{n}_{\alpha}$ (84)
into Eq. (82) and using Eq. (81), we derive
$\eta^{(n)}_{\alpha}=\lambda^{n}_{\alpha}-x\sum_{k=1}^{n}a_{k}\lambda^{n-k}_{\alpha},\quad a_{k}=\sum_{i=1}^{N}\epsilon_{i}^{k-1}\gamma_{i}^{2}=\langle\gamma|{\cal E}^{k-1}|\gamma\rangle,$ (85)
which also implies a recurrence relation for eigenvalues of $I_{n}$
$\eta^{(n+1)}_{\alpha}=\lambda_{\alpha}\eta^{(n)}_{\alpha}-xa_{n}.$ (86)
Note in particular that the eigenvalues of $H$ are
$\eta^{(1)}_{\alpha}=\lambda_{\alpha}-x\sum_{i}\gamma_{i}^{2}=\lambda_{\alpha}-x.$ (87)
The above expressions for the eigenvalues are helpful in establishing several facts. For example, $I_{n}$ is a polynomial in $H$ of order $n$. To write down this polynomial, observe that the eigenvalues of $H+x\mathds{1}={\cal E}+x{\bf\Gamma}$ are simply $\lambda_{\alpha}$, and Eqs. (85) and (86) therefore mean
$I_{n}=(H+x\mathds{1})^{n}-x\sum_{k=1}^{n}\langle\gamma|{\cal E}^{k-1}|\gamma\rangle(H+x\mathds{1})^{n-k},\quad I_{n+1}=HI_{n}+xI_{n}-x\langle\gamma|{\cal E}^{n}|\gamma\rangle.$ (88)
Further, replacing $H+x\mathds{1}$ with ${\cal E}+x{\bf\Gamma}$ in this equation and tracking down the terms linear in $x$ (higher orders in $x$ cancel identically), we derive $I_{n}$ explicitly in terms of ${\cal E}$ and ${\bf\Gamma}$,
$I_{n}={\cal E}^{n}+xK_{n}={\cal E}^{n}+x\sum_{k=1}^{n}\left({\cal E}^{n-k}{\bf\Gamma}{\cal E}^{k-1}-{\cal E}^{n-k}\langle\gamma|{\cal E}^{k-1}|\gamma\rangle\right).$ (89)
Finally, let us obtain the eigenvalues $\kappa^{(n)}_{\alpha}$ of $K_{n}$. Evidently,
$\kappa^{(n)}_{\alpha}=x^{-1}\lim_{x\to\infty}\eta^{(n)}_{\alpha}.$ (90)
As discussed above, in this limit $\lambda_{N}\to x\to\infty$. Remaining $\lambda_{\alpha}\to\tilde{\lambda}_{\alpha}$, where $\tilde{\lambda}_{\alpha}$ are finite and are the roots of the equation
$\sum_{i=1}^{N}\frac{\gamma_{i}^{2}}{\tilde{\lambda}_{\alpha}-\epsilon_{i}}=0,\quad\alpha=1,\dots,N-1.$ (91)
Plotting the left-hand-side of this equation as a function of $\tilde{\lambda}_{\alpha}$, we observe that the roots are sandwiched between $\epsilon_{i}$. Assuming $\epsilon_{i}$ are in ascending order, we have
$\epsilon_{1}<\tilde{\lambda}_{1}<\epsilon_{2}<\tilde{\lambda}_{2}<\dots<\epsilon_{N-1}<\tilde{\lambda}_{N-1}<\epsilon_{N}.$ (92)
Further, Eq. (82) implies $\kappa^{(n)}_{N}=0$ (these zero eigenvalues correspond to the common eigenvector $|\gamma\rangle$ of $K_{n}$), while Eq. (93) implies
$\kappa^{(n)}_{\alpha}=-\sum_{k=1}^{n}\langle\gamma|{\cal E}^{k-1}|\gamma\rangle\tilde{\lambda}^{m-k}_{\alpha},\quad\alpha=1,\dots,N-1.$ (93)
Suppose all eigenvalues $\epsilon_{i}$ of ${\cal E}$ are nonnegative and at least one of them is nonzero, i.e., ${\cal E}$ is a positive semi-definite matrix of nonzero rank. Then, Eqs. (92) and (93) imply that $\kappa^{(n)}_{\alpha}<0$ for $\alpha=1,\dots,N-1$. It follows that $|\gamma\rangle$ is the ground state of $-K_{n}$ with eigenvalue 0.
VI Quantum Computation with Type-1 Matrices
We now explore the potential application of Type-1 integrable matrices in quantum search algorithms. In the classical setting, locating a marked item in an unstructured database of size $N$ requires, on average, $N/2$ queries. Grover’s quantum algorithm Grover1 ; Grover2 offers a quadratic speedup, reducing the query complexity to $\mathcal{O}(\sqrt{N})$. We consider the adiabatic quantum computing implementation of this algorithm Farhi-Guttman ; QAA_Theory ; QAA ; LAA , in which the system evolves from a uniform superposition of all basis states to the target state under the specific time-dependent Hamiltonian $H_{G}(t)$ (referred to here as the Grover Hamiltonian). Remarkably, it turns out that $H_{G}(t)$ is Type-1 matrix—namely, the Hamiltonian defined in Eq. (53). We then demonstrate how the commuting operators $I_{n}$ can be exploited to enhance the fidelity of the quantum computation.
In quantum search algorithms Grover1 ; Grover2 ; Grover-Analysis ; QAA ; QAA_Theory ; Young ; Farhi-Guttman ; Childs ; Berkeley ; LAA , database items are represented as orthonormal basis states in a Hilbert space. The computation begins in the uniform superposition gamma
$|\gamma\rangle=\frac{1}{\sqrt{N}}\sum_{x=1}^{N}|x\rangle,$ (94)
which equally weights all possible configurations. The goal is to identify a specific, unknown target state $|m\rangle$. Grover’s algorithm achieves this by iteratively applying a sequence of unitary operations approximately $k\sim\pi\sqrt{N}/4$ times Grover1 ; Grover2 .
In contrast to this discrete-time protocol, we focus on an analog version of the quantum search Farhi-Guttman ; QAA_Theory ; QAA ; LAA . The system evolves under the time-dependent Grover Hamiltonian
$\begin{split}H_{G}(t)=s(t)\Bigl{[}\mathds{1}-|m\rangle\langle m|\Bigr{]}+\left[1-s(t)\right]\Bigl{[}\mathds{1}-|\gamma\rangle\langle\gamma|\Bigr{]},\end{split}$ (95)
with interpolation function $s(t)$ satisfying $s(0)=0$ and $s(T_{\mathrm{run}})=1$. In a perfectly adiabatic evolution, the system would follow the instantaneous ground state from $|\gamma\rangle$ at $t=0$ to the target state $|m\rangle$ at $t=T_{\mathrm{run}}$. Achieving this ideal evolution would require an infinitely long runtime, whereas the practical objective is to minimize $T_{\mathrm{run}}$ through a suitable choice of $s(t)$, while ensuring the final state remains sufficiently close to $|m\rangle$.
Roland and Cerf LAA showed that requiring the fidelity $F(t)$ between the evolving state $\Psi(t)$ and the instantaneous ground state $\Psi_{0}(t)$ to remain above a fixed threshold,
$F(t)=\left|\langle\Psi_{0}(t)|\Psi(t)\rangle\right|^{2}\geq F_{\min},$ (96)
throughout the evolution under the Hamiltonian (95)—that is, ensuring local adiabaticity—recovers the quadratic speedup of Grover’s algorithm. Assuming $F_{\min}$ is close to 1 and applying the adiabatic theorem, they further derive the corresponding interpolation function $s(t)$ as
$\begin{gathered}s(t)=\frac{N\tan\left(\frac{2t\delta}{N}\sqrt{N-1}\right)}{2\sqrt{N-1}\left[1+\sqrt{N-1}\tan\left(\frac{2t\delta}{N}\sqrt{N-1}\right)\right]},\end{gathered}$ (97)
where $\delta=\sqrt{1-F_{\min}}$. The complexity of this locally adiabatic evolution governed by $H_{G}(t)$ is then determined by the condition $s(T_{\mathrm{run}})=1$, which yields
$T_{\mathrm{run}}=\frac{1}{\delta}\frac{N}{\sqrt{N-1}}\arctan{\sqrt{N-1}}\;\;\;\stackrel{{\scriptstyle\mathclap{\tiny\mbox{$N\gg 1$}}}}{{\approx}}\;\;\;\frac{\pi}{2\delta}\sqrt{N}.$ (98)
We observe that the Grover Hamiltonian $H_{G}(t)$ is a special case of the Type-1 Hamiltonian $H$. To see this, note that the Hamiltonian $H$ defined in Eq. (53), along with its commuting partners $I_{n}$ in Eq. (56), can be viewed as a two-parameter family of operators,
$H=u\,{\cal E}+v\,(\mathds{1}-{\bf\Gamma}),\quad I_{n}=u\,{\cal E}^{n}-v\,K_{n},$ (99)
where we have redefined $x=-v/u$ and rescaled all operators by an overall factor of $u$. This reparameterization leaves the commutation relations $[H,I_{n}]=[I_{k},I_{n}]=0$ unchanged. We now fix the basis in which ${\cal E}$ is diagonal so that $|1\rangle=|m\rangle$, and choose the parameters
$\epsilon_{1}=0,\quad\epsilon_{2}=\cdots=\epsilon_{N}=1.$ (100)
With this choice, ${\cal E}=\mathds{1}-|m\rangle\langle m|$, and since ${\bf\Gamma}=|\gamma\rangle\langle\gamma|$ by definition, we find that $H=H_{G}(t)$ along the one-parameter trajectory $u=s(t)$, $v=1-s(t)$ in the $(u,v)$ parameter space. The fact that $H_{G}(t)$ lies within the Type-1 family is notable, given that Type-1 matrices form a measure-zero subset of all matrices of the form $M(u,v)=uA+vB$, where $A$ and $B$ are arbitrary real symmetric matrices independent of $u$ and $v$ haile ; haile1 .
Figure 1: Efficiency increase of the Grover’s search algorithm by utilizing quantum integrability of Type-1 matrices. Panel (a): Instantaneous energy gaps between the ground and first excited state as functions of $t\in(0,T_{\mathrm{run}})$ for two random sets of $\epsilon_{i>2}$ with $\Delta\epsilon=0.1$ and three different currents $I_{n}$ as well as for the Grover Hamiltonian (solid black line). Panel (b): $(1-F_{n})$ for different $I_{n}(t)$ and the same random sets of $\epsilon_{i>2}$ as in panel (a), where $F_{n}$ is the fidelity. The increased fidelity as compared to the Grover Hamiltonian (horizontal black line) is due to quantum interference. This effect is particularly dramatic for $n=3$ and $n=7$ with $1-F_{3}=5.20\times 10^{-6}$ and $1-F_{7}=6.40\times 10^{-5}$ for the random choice #2 RandSeed . The matrix size is $N=64$ and the parameter $\delta=0.1$ in the interpolating function $s(t)$.
As established in the previous section, when ${\cal E}$ is positive semi-definite with nonzero rank, all operators $-K_{n}$ share the same ground state $|\gamma\rangle$. Likewise, when $\epsilon_{1}=0$ and all other $\epsilon_{i}>0$, the operators ${\cal E}^{n}$ share the common ground state $|1\rangle=|m\rangle$. It follows that the same adiabatic quantum computation can be implemented using any of the commuting $I_{n}(t)$ as the interpolating Hamiltonian,
$I_{n}(t)=s(t)\,{\cal E}^{n}+\left[1-s(t)\right](-K_{n}).$ (101)
This raises a natural question: how does the choice of $I_{n}(t)$ affect the accuracy of the computation? To address this, we compare the fidelities
$F_{n}=\left|\left\langle m\middle|\Psi_{(n)}(T_{\mathrm{run}})\right\rangle\right|^{2},$ (102)
by numerically evaluating $\Psi_{(n)}(t)$, the solution of the non-stationary Schrödinger equation $i\,\partial_{t}\Psi=I_{n}(t)\,\Psi$, with the initial condition $|\Psi(0)\rangle=|\gamma\rangle$. For simplicity, we use the same interpolating function $s(t)$ given by Eq. (97) for all $I_{n}(t)$.
With the choice of $\epsilon_{i}$ in Eq. (100), all operators $I_{n}$, including $I_{1}\equiv H=H_{G}$, are block-diagonal, consisting of a common $2\times 2$ block and an $(N-2)\times(N-2)$ block proportional to the identity matrix. This $2\times 2$ block governs the time evolution of the initial uniform superposition state $|\gamma\rangle$ under $I_{n}$, and hence the fidelities $F_{n}$ are identical for all $n$. To see this, note that
$|\gamma\rangle=\frac{1}{\sqrt{N}}|1\rangle+\frac{\sqrt{N-1}}{\sqrt{N}}\,|\tilde{2}\rangle,$ (103)
[…]
According to Eqs. (82) and (81), the eigenvalues $\eta^{(n)}_{1}$ and $\eta^{(n)}_{2}$ of the $2\times 2$ block of $I_{n}$, for the choice of $\epsilon_{i}$ in Eq. (100) and $\gamma_{i}=1/\sqrt{N}$, are given by
$\eta^{(n)}_{1,2}=x\frac{N-1}{N}\cdot\frac{1}{\lambda_{1,2}-1},$ (104)
where $\lambda_{1}$ and $\lambda_{2}$ are the roots of the quadratic equation
$\frac{1}{N}\left(\frac{1}{\lambda}+\frac{N-1}{\lambda-1}\right)=\frac{1}{x}.$ (105)
Since the $I_{n}$ commute, their $2\times 2$ blocks must commute as well. Commuting $2\times 2$ matrices with identical eigenvalues must coincide.
Figure 2: Efficiency gain using $I_{3}$ for matrix size $N=64$. Panel (a): Five values of $\Delta\epsilon$, spaced logarithmically, are tested with fixed $\delta=0.1$. Random choices #$\#1$ and #$\#2$ are shown as blue crosses and red squares, respectively RandSeed . As $\Delta\epsilon\to 0$, fidelity-deficit converges to that of the Grover Hamiltonian (black line). Notably, the performance of $I_{3}$ improves markedly at $\Delta\epsilon=0.1$. Panel (b): Holding $\Delta\epsilon=0.1$ fixed, we vary $\delta$ using the same random $\epsilon_{i}$ as in Fig. 1. As expected, both $T_{\mathrm{run}}$ and $F_{3}$ rise with decreasing $\delta$. Curiously, lowering $\delta$ brings out a marked efficiency gain in random choice #$\#1$, driven by quantum integrability.
We now turn to other choices of positive semi-definite ${\cal E}$ with $\epsilon_{1}=0$ and all other $\epsilon_{i}>0$. As shown above, the same quantum search protocol can be implemented for any such ${\cal E}$ using any member $I_{n}$ of the commuting Type-1 family. Consider, for example, the following parameter choice:
$\begin{split}&\epsilon_{1}=0,\quad\epsilon_{2}=1,\\ &\epsilon_{i}=\text{uniformly distributed random numbers in the interval }[1.0,1.0+\Delta\epsilon],\quad i>2,\end{split}$ (106)
where $\Delta\epsilon$ is a small positive number. By construction, the energy gap between the ground and first excited states at $t=T_{\mathrm{run}}$ is the same (equal to 1) for all $I_{n}(t)$ and for the Grover Hamiltonian (95). In addition, this gap is also fixed at 1 for all $I_{1}(t)\equiv H(t)$ at $t=0$, regardless of the specific choice of $\epsilon_{i}$ for $i>2$. In this case, the Hamiltonian and the operators $I_{n}$ are no longer block-diagonal. Consequently, the instantaneous ground state is coupled not only to the first excited state but also to higher excited states, and the fidelities $F_{n}$ become distinct. As shown in Fig. 1 panel (a), the gap between the instantaneous ground state and the first excited state is nearly the same as that for the Grover Hamiltonian (95) across all $I_{n}$.
At first glance, one might expect that quantum tunneling to higher excited states would reduce the fidelity by increasing probability leakage out of the ground state. Surprisingly, our numerical simulations indicate the opposite behavior; see Fig. 1. The fidelity $F_{n}$ exceeds that of the Grover Hamiltonian (95) for all $I_{n}$, and for $n=3$, fidelity-deficit $1-F_{n}$ is smaller by two orders of magnitude. This enhancement is a result of quantum interference: probability amplitudes for transitions out of the ground state combine destructively, suppressing leakage into excited states.
Fig. 2 examines $F_{3}$ more closely for $n=3$, $N=64$. In panel (a), two sets of random $\epsilon_{i>2}$ are drawn for various $\Delta\epsilon$, keeping $\delta=0.1$ fixed. As $\Delta\epsilon\to 0$, the fidelity approaches that of the Grover Hamiltonian. In this instance, the best performance occurs for random choice #$\#2$ at $\Delta\epsilon=0.1$. Panel (b) explores the effect of varying $\delta$ at fixed $\Delta\epsilon=0.1$. As expected, lowering $\delta$ improves fidelity — at the cost of a longer $T_{\mathrm{run}}$.
It is worth noting that Type-1 matrices also give rise to certain integrable multi-level Landau-Zener models Patra_2015 ; Yuzbashyan_2018 . These models are known to exhibit similarly favorable quantum interference effects among transition amplitudes demkov ; Sinitsyn_2015 ; Sinitsyn_2017 . Our results thus suggest that leveraging the integrability of Type-1 matrices in adiabatic quantum computation can lead to significant improvements in performance.
VII Concluding Remarks
In this work, we have noted that Weyl’s relations involving two unitary matrices $A$ and $B$ in $N$ dimensions can be usefully generalized to include a third matrix $C$. The matrix $C$ has interesting commutation relations with $\log A$ which permits one to view it as a finite dimensional version of the operator $\hat{Q}\hat{P}+\hat{P}\hat{Q}$ in quantum mechanics, while its commutation relation with $B$ enables us to view it as the canonical momentum $\hat{P}$ when restricted to an $(N-1)$-dimensional subspace obtained by eliminating a specific state of the original state space. A simple procedure for modifying arbitrary operators (matrices) to annihilate the specific state then generates an effectively finite-dimensional canonical theory. These matrices are then used to construct a hierarchy of mutually commuting operators, depending on $\sim 2N$ independent real parameters. These operators are shown to be intimately connected to Type-1 matrices developed by us earlier yuzbashyan , using very different and independently developed ideas. The Type-1 matrices provide a realization of quantum integrable systems—such as the Heisenberg antiferromagnetic spin chain, or the one-dimensional Hubbard model, so that the new hierarchy can be viewed in the same light.
Additionally, we find that the hierarchy of commuting operators obtained here contain the leading member $I_{1}$, given by Eq. (53), which, for a certain choice of the parameters, is identical to the Grover Hamiltonian (95), widely studied in the field of quantum computation. We present results that suggest a significant advantage of employing higher members $I_{n}$ with $n\geq 2$ in similar studies, which seem to be worth pursuing further.
Finally, we find it most remarkable that Weyl’s ideas Weyl continue to nourish current research and open new directions, nearly a century after their original publication.
Acknowledgements.
We are grateful to Dominik Šafránek for bringing Ref. LAA to our attention. AP acknowledges the financial support from the Institute for Basic Science (IBS) in the Republic of Korea through Project No. IBS-R024-D1.